OpenStudy (anonymous):

a coin tossed 60 times about how many times would you expect it to land on heads

OpenStudy (muzzack):

well if the probability of having heads is 1/2 then multiply 1/2 and 60

OpenStudy (muzzack):

30

OpenStudy (muzzack):

so it would be \[\frac{ 1 }{ 2 } \times 60 = 30\]

OpenStudy (ybarrap):

Using the Binomial Distribution, the average number of successes is \(n\times p\), where \(n=60\) and \(p=\cfrac{1}{2}\): http://en.wikipedia.org/wiki/Binomial_distribution#Mean_and_variance